OpenStudy (patrickweewoos):

A hot tub holding 480 gal of water begins to empty at a rate of 5 gal/min. At the same time, an empty hot tub begins to fill at a rate of 15 gal/min. Let g represent the number of gallons of water and let t represent time in minutes. The system models this situation. How long will it take for the hot tubs to hold equal amounts of water, and how much water will that be? g = 480 – 5t g = 15t It will take______ minutes for both hot tubs to hold equal amounts of water. They will each hold_____gallons.

OpenStudy (patrickweewoos):

@perl

OpenStudy (perl):

solve 480 - 5t = 15t 480 = 20t t = 24 minutes # gallons = 360

OpenStudy (patrickweewoos):

What?

OpenStudy (patrickweewoos):

@perl

OpenStudy (patrickweewoos):

24 minuets and 360 gallons

OpenStudy (patrickweewoos):

@perl

OpenStudy (perl):

yes

OpenStudy (patrickweewoos):

OK :D, Farrah needs an electrician and must decide between 2 companies. For a service visit, Company A charges $50 to send an electrician plus $40/h. Company B charges $60 to send an electrician plus $36/h. The graph that models this situation is shown here. According to the graph, how long must each electrician work in order for the two to charge equal amounts? hour(s)

OpenStudy (patrickweewoos):

Thats the graph

OpenStudy (patrickweewoos):

@perl Help plz

OpenStudy (patrickweewoos):

Did you see the graph

OpenStudy (perl):

yes 2.5 hours

OpenStudy (patrickweewoos):

ok now @perl , Paul’s Pool Service charges $340 to open a pool and charges $28 each week for pool maintenance. Nelson’s Pool Service charges $280 to open a pool and charges $32 each week for pool maintenance. The system that models this situation is given, where c is the cost of pool maintenance and w is the number of weeks. c = 340 + 28w c = 280 + 32w The solution to the system is (15, 760). Which interpretation correctly describes the solution to the system of equations? A. At 15 wk of pool cleanings, a customer will reach the maximum payment of $535. B. At 15 wk, the cost for pool maintenance is the same for both companies. Both companies will charge $760. C. Paul’s Pool Service charges more money in the fifteenth week by charging $760. D. Nelson’s Pool Service charges more money in the fifteenth week by charging $760.

OpenStudy (patrickweewoos):

@perl

OpenStudy (patrickweewoos):

@perl

OpenStudy (patrickweewoos):

@perl

OpenStudy (patrickweewoos):

@iGreen

OpenStudy (patrickweewoos):

@ganeshie8

OpenStudy (patrickweewoos):

@perl only 2 more

OpenStudy (anonymous):

lol